OpenStudy (anonymous):

Please Help! MEDAL WILL BE REWARDED!!! Find d^2(y)/dx^2 in terms of x and y. y^2=4x

OpenStudy (jjuden):

It's just asking you to find the second derivative using implicit differentiation. y^2 = 4x 2yy' = 4 yy' = 2 y(y'') + y'(y') = 0 Substitute in for y and y' sqrt(4x)y'' + (2/y)^2 = 0 y'' = -(2/y)^2 / sqrt(4x)

OpenStudy (anonymous):

okay, so that is in terms of both x and y?

OpenStudy (anonymous):

Where does the sqroot come from?

OpenStudy (jjuden):

it was in the origanal problem

OpenStudy (anonymous):

oh, I see okay, thank you!

OpenStudy (jjuden):

welcome hope i helped some

OpenStudy (anonymous):

Is your last answer my final answer or do I simplify? @jjuden

OpenStudy (jjuden):

you have to Substitute in for y and y'